# The effects of detuning on entropic uncertainty bound and quantum correlations in dissipative environment

### 19 Jan 2024 v2

@@ 1 Introduction
where $\Pi_{x}^{A}=|x_{i}\rangle\langle x_{i}|\otimes\mathrm{I}$ and $\mathrm{I}$ is the identity operator. Numerous studies have been conducted about the entropic uncertainty relation with quantum memory EUR-QM [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65] . So far, Adabi et al. have been proposed the tightest bound as follow [61]
@@ 2 Dynamics of two-level open systems in the presence of detuning
where $\delta=\omega_{0}-\omega$ is the detuning between the transition frequency of the atom $\omega_{0}$ and the center frequency of the cavity $\omega_{c}$. As a result of detuning, the effective coupling between a qubit and the environment is decreased. For Lorentzian spectral density, the spectral width of the environment shows by $\lambda$ which associated with the correlation time of the environment as $\tau_{c}=\lambda^{-1}$. The time scale during which the state of the atom changes depends on parameter $\gamma$ through $\tau_{a}=\gamma^{-1}$ [109]. By applying the Lorentzian spectral density, the correlation function of the environment $f(t-t_{1})$ is obtained as follows
@@
$G(t)=e^{-(\lambda-i\delta)t/2}\left[\cosh\left(\frac{\Omega t}{2}\right)-\frac{\lambda-i\delta}{\Omega}\sinh\left(\frac{\Omega t}{2}\right)\right],$ (19)
@@
$\frac{\partial}{\partial t}\rho(t)=\Gamma(t)\left(\sigma_{\rho}(t)\sigma_{+}-\frac{1}{2}\left\{\sigma_{+}\sigma_{-},\rho(t)\right\}\right).$ (20)
@@ 3 Quantum correlations and entropic uncertainty bound with detuning
In this section, the effects of detuning on quantum correlations and entropic uncertainty bound will be studied. Here, we are considering the scenario in which the quantum memory $B$ interacts with the environment while the measured particle $A$ remains unaffected by environment and Alice does her measurement on it. The dynapics of bipartite state $\rho^{AB}$ can be described as
@@
Let us consider the case in which the following two-qubit state have been shared between Alice and Bob as
@@
parameter $\theta$ is considered the correlation parameter. In Eq.26 first qubit is considered as measured particle and the sconde one is considered a quantum memory. It can be , QD has its maximum value one for $\theta=\pi/4$ and $43\pi/$ 4 with $r=1$.
@@
$\rho_{14}$ $=$ $\rho_{41}^{\star}=G(t)\cos\theta\sin\theta,$ (29)
@@ 3.1 Entanglement and Quantum discord
Several criteria are used in the measurement of quantum correlations, as mentioned in the introduction. In this study, we measure quantum correlations using two practical and optimal criteria: concurrence(entanglement measyre) and QD. The concurrence for two-qubit states can be defined as [113]
@@
where $\varepsilon_{i}$’s are the eigenvalues of $R=\sqrt{\sqrt{\rho}\tilde{\rho}\sqrt{\rho}}$, in decreasing order and $\tilde{\rho}=(\sigma_{y}\otimes\sigma_{y})\rho^{\star}(\sigma_{y}\otimes\sigma_{y})$ with $\sigma_{y}$ is the $y$ compinent of Pauli matrices and $\rho^{\star}$ is the complex conjugate of $\rho$. The concurrence of the state in Eq.28 can be obtained as
$\mathcal{C}(\rho)=2\max\{0,|\rho_{14}|-\sqrt{\rho_{22}-\rho_{33}}\}.$ (31)
@@
$\mathcal{Q}_{j}=h(\rho_{11}+\rho_{33})+\sum_{j=1}^{4}\lambda_{i}\log_{2}\lambda_{i}+\mathcal{D}_{j}$ (33)
and
@@
$\mathcal{D}_{2}=-\sum_{i}\rho_{ii}\log_{2}\rho_{ii}-\mathcal{H}\left(\rho_{11}+\rho_{33}\right),$
@@ 3.2 Entropic uncertainty bound
$I\left(\sigma_{z};B\right)=$ $-\left(\rho_{11}+\rho_{22}\right)\log_{2}\left(\rho_{11}+\rho_{22}\right)$ (35)
$-\left(\rho_{33}+\rho_{44}\right)\log_{2}\left(\rho_{33}+\rho_{44}\right)$
$-\left(\rho_{11}+\rho_{33}\right)\log_{2}\left(\rho_{11}+\rho_{33}\right)$
$-\left(\rho_{22}+\rho_{44}\right)\log_{2}\left(\rho_{22}+\rho_{44}\right)$
$+\sum_{i}\rho_{ii}\log_{2}\rho_{ii},$
@@
$I\left(\sigma_{x};B\right)=$ $-\left(\rho_{11}+\rho_{33}\right)\log_{2}\left(\rho_{11}+\rho_{33}\right)$ (36)
$-\left(\rho_{22}+\rho_{44}\right)\log_{2}\left(\rho_{22}+\rho_{44}\right)$
$+\sum_{i}\xi_{i}\log_{2}\xi_{i}+1$
where $\xi_{1}=\xi_{2}=\frac{1-k}{4}$ and $\xi_{3}=\xi_{4}=\frac{1+k}{4}$ with $k=\sqrt{4\left|\rho_{14}+\right|^{2}+\left[1-2\left(\rho_{22}+\rho_{44}\right)\right]^{2}}$. From Eq. 28, the mutual information $\mathcal{I}(\rho_{AB}(t))$ can be obtained as
$\mathcal{I}(A;B)=$ $-\left(\rho_{11}+\rho_{22}\right)\log_{2}\left(\rho_{11}+\rho_{22}\right)$ (37)
$-\left(\rho_{33}+\rho_{44}\right)\log_{2}\left(\rho_{33}+\rho_{44}\right)$
$-\left(\rho_{11}+\rho_{33}\right)\log_{2}\left(\rho_{11}+\rho_{33}\right)$
$-\left(\rho_{22}+\rho_{44}\right)\log_{2}\left(\rho_{22}+\rho_{44}\right)$
$+\sum_{i}\lambda_{i}\log_{2}\lambda_{i},$
@@ 4 Results
In this section we study the behavior of the EUB and quantum correlation in the presence of detuning in dissipative environment. We consider the maximally entangled state with $r=1$, $\theta=\pi/4$ as the initial state which has been shared between ALice and Bob. By setting $r=1$ and $\theta=\pi/4$ in Eq. 26, we have the maximally entangled initial state which has been shared between Alice and Bob. In Fig.2, quantum correlations has been plotted as a function of dimensionless parameter $\gamma t$ in Markovian regime with $\lambda=10\gamma$ for different values of detuning $\delta$. Fig.2.(a), shows the QD in Markovian regime for different values of detuning $\delta$. As can be seen increasing $\delta$ protects the QD from decay. In other words, increasing detuning parameter leads to more preservation of QD in the presence of the dissipative environment. In Fig.2(b), concurrence has been plotted as a function of dimensionless parameter $\gamma t$. As can be seen the same results have been obtained for concurrence. So, it can be concluded that increasing detuning can protect quantum correlation from decoherence.
@@
In Fig.4, quantum correlations includes entanglement and QD have been plotted as a function of dimensionless parameter $\gamma t$ for different values of detuning parameter in non-Markovian regime with $\lambda=0.1\gamma$ for maximally entangled initial state. Fig.4(a), shows QD in terms of dimensionless for different values of detuning parameter in non-Markovian regime. As can be seen, can be preserve during the evolution by increasing detuning parameter. In Fig.4(b), the concurrence as a measure for entanglement has been plotted as a function of detuning parameter in non-Markovian regime for different values of detuning parameter. Again, one can see the concurrence can be preserve during interaction with dissipative environment by increasing the detuning parameter.
@@
In Fig.5, the EUB has been plotted as a function of dimensionless parameter $\gamma t$ in non-Markovian environment for different values of detuning parameters $\delta$. As expected, it can be seen that the EUB decrases with increasing the detuning parameter.
@@ 5 Conclusion
In this work, the effects of detuning on quantum correlations and entropic uncertainty bound have been studied. It was observed that there exist inverse relation between entopic uncertainty bound and quantum correlation between measured particle and quantum memory. It was also shown that the quantum correlation increases with increasing detuning between the transition frequency of a quantum memory and the center frequency of a cavity. The situation for entropic uncertainty bound was quite different. It was observed that the entropic uncertainty bound can be reduced with increasing the detuning parameter. From these results, it can be concluded that the correlation between the measured particle, which is in the possession of Alice and quantum memory , which is in the possession of Bob, has an important role in guessing the result of Alice’s measurement by Bob.

### 25 Jan 2024 v3

@@ 1 Introduction
where $\Pi_{x}^{A}=|x_{i}\rangle\langle x_{i}|\otimes\mathrm{I}$ and $\mathrm{I}$ is the identity operator. Numerous studies have been conducted about the entropic uncertainty relation with quantum memory EUR-QM [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76] . So far, Adabi et al. have been proposed the tightest bound as follow [72]
@@ 2 Dynamics of two-level open systems in the presence of detuning
where $\delta=\omega_{0}-\omega_{c}$ is the detuning between the transition frequency of the atom $\omega_{0}$ and the center frequency of the cavity $\omega_{c}$. As a result of detuning, the effective coupling between a qubit and the environment is decreased. For Lorentzian spectral density, the spectral width of the environment shows by $\lambda$ which associated with the correlation time of the environment as $\tau_{c}=\lambda^{-1}$. The time scale during which the state of the atom changes depends on parameter $\gamma$ through $\tau_{a}=\gamma^{-1}$ [120]. By applying the Lorentzian spectral density, the correlation function of the environment $f(t-t_{1})$ is obtained as follows
@@
$G(t)=e^{-(\lambda-i\delta)t/2}\left[\cosh\left(\frac{\Omega t}{2}\right)+\frac{\lambda-i\delta}{\Omega}\sinh\left(\frac{\Omega t}{2}\right)\right],$ (19)
@@
$\frac{\partial}{\partial t}\rho(t)=\Gamma(t)\left(\sigma_{-}\rho(t)\sigma_{+}-\frac{1}{2}\left\{\sigma_{+}\sigma_{-},\rho(t)\right\}\right).$ (20)
@@ 3 Quantum correlations and entropic uncertainty bound with detuning
In this section, the effects of detuning on quantum correlations and entropic uncertainty bound will be studied. Here, we are considering the scenario in which the quantum memory $B$ interacts with the environment while the measured particle $A$ remains unaffected by environment and Alice does her measurement on it. The dynamics of bipartite state $\rho^{AB}$ can be described as
@@
Let us consider the case in which the following two-qubit state has been shared between Alice and Bob as
@@
parameter $\theta$ is considered the correlation parameter. In Eq.26 first qubit is considered as measured particle and the second one is considered a quantum memory. QD has its maximum value one for $\theta=\pi/4$ and $43\pi/$ 4 with $r=1$.
@@
$\rho_{14}$ $=$ $\rho_{41}^{\star}=G(t)r\cos\theta\sin\theta,$ (29)
@@ 3.1 Entanglement and Quantum discord
Several criteria are used in the measurement of quantum correlations, as mentioned in the introduction. In this study, we measure quantum correlations using two practical and optimal criteria: concurrence(entanglement measure) and QD. The concurrence for two-qubit states can be defined as [124]
@@
where $\varepsilon_{i}$’s are the eigenvalues of $R=\sqrt{\sqrt{\rho}\tilde{\rho}\sqrt{\rho}}$, in decreasing order and $\tilde{\rho}=(\sigma_{y}\otimes\sigma_{y})\rho^{\star}(\sigma_{y}\otimes\sigma_{y})$ with $\sigma_{y}$ is the $y$ component of Pauli matrices and $\rho^{\star}$ is the complex conjugate of $\rho$. The concurrence of the state in Eq.28 can be obtained as
$\mathcal{C}(\rho)=2\max\{0,|\rho_{14}|-\sqrt{\rho_{22}\rho_{33}}\}.$ (31)
@@
$\mathcal{Q}_{j}=h(\rho_{11}+\rho_{33})+\sum_{i=1}^{4}\lambda_{i}\log_{2}\lambda_{i}+\mathcal{D}_{j}$ (33)
where $j\in\{1,2\}$ and
@@
$\mathcal{D}_{2}=-\sum_{i}\rho_{ii}\log_{2}\rho_{ii}-h\left(\rho_{11}+\rho_{33}\right),$
@@ 3.2 Entropic uncertainty bound
$I\left(\sigma_{z};B\right)=h(\rho_{11}+\rho_{22})+h(\rho_{11}+\rho_{33})+\sum_{i}\rho_{ii}\log_{2}\rho_{ii},$ (35)
@@
$I\left(\sigma_{x};B\right)=1+h(\rho_{11}+\rho_{33})+\sum_{i}\xi_{i}\log_{2}\xi_{i},$ (36)
where $\xi_{1}=\xi_{2}=\frac{1-k}{4}$ and $\xi_{3}=\xi_{4}=\frac{1+k}{4}$ with $k=\sqrt{4\left|\rho_{14}\right|^{2}+\left[1-2\left(\rho_{22}+\rho_{44}\right)\right]^{2}}$. From Eq. 28, the mutual information $\mathcal{I}(\rho_{AB}(t))$ can be obtained as
$I(A;B)=h(\rho_{11}+\rho_{22})+h(\rho_{11}+\rho_{33})+\sum_{i}\lambda_{i}\log_{2}\lambda_{i},$ (37)
@@ 4 Results
In this section we study the behavior of the EUB and quantum correlation in the presence of detuning in dissipative environment. We consider the maximally entangled state with $r=1$, $\theta=\pi/4$ as the initial state which has been shared between ALice and Bob. By setting $r=1$ and $\theta=\pi/4$ in Eq. 26, we have the maximally entangled initial state which has been shared between Alice and Bob. In Fig.2, quantum correlations has been plotted as a function of dimensionless parameter $\gamma t$ in Markovian regime with $\lambda=10\gamma$ for different values of detuning $\delta$. Fig.2.(a), shows the QD in Markovian regime for different values of detuning $\delta$. It can be seen that increasing $\delta$ can protect the QD from decay. In other words, increasing detuning parameter leads to more preservation of QD in the presence of the dissipative environment. In Fig.2(b), concurrence has been plotted as a function of dimensionless parameter $\gamma t$. As can be seen the same results have been obtained for concurrence. So, it can be concluded that increasing detuning can protect quantum correlation from decoherence.
@@
In Fig.4, entanglement and QD have been plotted as a function of dimensionless parameter $\gamma t$ for different values of detuning parameter in non-Markovian regime with $\lambda=0.1\gamma$ for maximally entangled initial state. Fig.4(a), shows QD in terms of dimensionless for different values of detuning parameter in non-Markovian regime. As can be seen, QD can be protected from decoherence during the evolution by increasing detuning parameter. In Fig.4(b), the concurrence as a measure for entanglement has been plotted as a function of detuning parameter in non-Markovian regime for different values of detuning parameter. Again, one can see the concurrence can be preserve during interaction with dissipative environment by increasing the detuning parameter.
@@
In Fig.5, the EUB has been plotted as a function of dimensionless parameter $\gamma t$ in non-Markovian environment for different values of detuning parameters $\delta$. As expected, it can be seen that the EUB decreases with increasing the detuning parameter.
@@ 5 Conclusion
In this work, the effects of detuning on quantum correlations and entropic uncertainty bound have been studied. It was observed that there exist inverse relation between entropic uncertainty bound and quantum correlation between measured particle and quantum memory. It was also shown that the quantum correlation increases with increasing detuning between the transition frequency of a quantum memory and the center frequency of a cavity. The situation for entropic uncertainty bound was quite different. It was observed that the entropic uncertainty bound can be reduced with increasing the detuning parameter. From these results, it can be concluded that the correlation between the measured particle, which is in the possession of Alice and quantum memory , which is in the possession of Bob, has an important role in guessing the result of Alice’s measurement by Bob.